# arXiv:2511.09500v3[stat.ML] 01 Mar 2026

### v3

@@ 5.1 First-Order Denoising Equation
Theorem 3 (Distributional Matching: Optimal First-Order Denoiser).
Under Assumptions 4(i) and 5(i), we consider the first-order denoiser $\mathbf{T}_{1}^{f}(y)$.
Let $p=P_{X}$ satisfy Assumption 1 for up to $k=4$, and the noise $Z$ has bounded 4-th moments as in Assumption 2(i). Then with some universal constant $C_{\Omega}>0$,
@@
An immediate corollary is a lower bound showing that the classical Bayes-optimal denoiser only achieves $\Theta(\sigma^{2})$ accuracy in terms of distributional matching, thus is sub-optimal compared to the first-order denoiser $\mathbf{T}_{1}^{f_{*}}$, in the small constant noise regime.
Corollary 2 (Distributional Matching: Bayes-optimal Denoiser).
Consider the same setting and assumptions as in Theorem 3. The Bayes-optimal denoiser
@@ 5.2 Second-Order Denoising Equation
Theorem 4 (Distributional Matching: Optimal Second-Order Denoiser).
Under Assumptions 4(i) & (ii) and 5 (i) & (ii), we consider the second-order denoiser $\mathbf{T}_{2}^{f,g}$.
Let $p=P_{X}$ satisfy Assumption 1 for up to $k=6$, and the noise $Z$ has bounded 6-th moments as in Assumption 2(i) & (ii). Then with some universal constant $C_{\Omega}>0$,
@@
Again, this theorem requires stronger conditions than Theorem 3, in terms of smoothness of the signal distribution $P_{X}$ (up to 6-th order), and boundedness of 6-th moments and a Gaussian-like 4-th moment assumption, stated in Assumption 2(ii). If instead one wishes to consider general noise $Z$ with i.i.d. coordinates that have general 4-th moments $\mathop{\mathbb{E}}[Z_{i}^{4}]=\kappa\in[1,\infty)$, then the second-order denoising equation will be modified by adding a term of the form $(3-\kappa)\partial_{i}^{3}q$ for each coordinate $i$ in (5.2). We choose to state a less general version for simplicity of presentation.
@@ 6 Numerics
In both experiments, we observe the following phenomenon. First, the Bayes-optimal denoiser introduces an overly aggressive shrinkage, thus shrinking the distribution $\mathbf{T}^{\star}\sharp P_{Y}$ inside the mass of $P_{X}$ and producing poor distributional matching. In the mixture of Gaussian experiment, the $W(\mathbf{T}^{\star}\sharp P_{Y},P_{X})=0.214$, not a good improvement compared to no shrinkage $W(P_{Y},P_{X})=0.233$. In contrast, both the first-order and second-order denoisers produce significantly better distributional matching, with $W(\mathbf{T}_{1}^{f_{*}}\sharp P_{Y},P_{X})=0.050$ and $W(\mathbf{T}_{2}^{f_{*},g_{*}}\sharp P_{Y},P_{X})=0.038$, an order of magnitude improvement. Second, although visually the first-order and second-order denoisers look similar, the second-order denoiser achieves a better Wasserstein distance in both experiments by a margin, demonstrating the benefit of higher-order denoising. Both of our proposed denoisers significantly outperform the Bayes-optimal denoiser in terms of distributional matching, both visually and quantitatively. This is compatible with our theoretical findings in Theorems 1-4.
@@ A.1 Proofs for Section 4
[…]
@@
[…]
@@
[…]
Moving to the first term on the right-hand side of (A.2), by Lemma 1(i), we have
@@
[…]
if (i) $\int_{\mathbb{R}^{d}}|m(y)|\differential y<\infty, d$$\int_{\mathbb{R}^{d}}|\partial_{i}m(y)|\differential y<\infty$, and (ii) $m\nabla p$ vanishes as $y\rightarrow\infty$ because $m$ vanishes and $\nabla p$ is bounded due to Assumption 1. To be more specific, (ii) is a direct application of integration by parts
@@
[…]
@@
[…]
@@
Proof of corollary 1.
Recall the last equation in the proof of Theorem 1,
[…]
@@
Proof of Theorems 2.
Again, let $\eta:=\sigma^{2}/2$ to streamline the calculations. For simplicity, we use $f=f_{*}$ and $g=g_{*}$, defined in (3.2), in the following proof whenever there is no ambiguity.
@@
[…]
@@
The constant in the term $O(\eta^{3})$ depends on $\sup_{y\in\mathbb{R}^{d}}\|\nabla^{3}m(y)\|_{\infty}$ and $\|\nabla h(y)\|^{3}$, with the first bounded (Assumption 1) and second integrable (Assumption 3).
@@
[…]
@@
and the boundary condition as $m\nabla p,m\nabla\Delta p$ vanishes on the boundary as $y\rightarrow\infty$. Here we use the fact that $m$ vanishes and $\nabla p,\nabla\Delta p$ are bounded (Assumption 1). Now invoke Lemma 1(ii) again, we have element-wise the following expansions
@@
[…]
@@
Note that $q\langle\nabla m,\nabla f\rangle\nabla f$ vanishes on the boundary as $y\rightarrow\infty$ due to the fact that $\partial_{i}m$ vanishes and $q\partial_{i}f\partial_{j}f$ is bounded (Assumption 3). Therefore, applying integration by parts and recalling (A.7), we have
[…]
=d[$=-\int_{\mathbb{R}^{d}}\left[\langle\nabla m(y),\nabla^{2}f(y)\nabla f(y)\rangle q(y)+\langle\nabla m(y),\nabla f(y)\rangle\langle\nabla f(y),\nabla q(y)\rangle+\langle\nabla m(y),\Delta f(y)\nabla f(y)\rangle q(y)\right]\differential{y}\;.$
@@
[…]
Using Lemma 2, and recall that $f_{*}=\log q$, we have
@@ A.2 Proofs for Section 5
Proof of Theorem 3.
Again, recall the scaling $\eta:=\sigma^{2}/2$, and set $f=f^{\star}$ for simplicity of notation. By definition of $d_{\mathrm{MA}}$, and that $p=P_{X}$, $q=P_{Y}$, we have
@@
where $O(\eta^{2})$ is uniform in $y$ under Assumption 1 for up to $k=4$.
@@
Here the $O(\eta^{2})$ term is uniform in $y$ under Assumption 1 for up to $k=2$ and Assumption 5(i).
Moreover, by the expansion of the determinant, and Assumption 5(i), we have
@@
Again, the $O(\eta^{2})$ term is uniform in $y$ ensured by Assumption 5(i).
@@
Together with the assumption that $\nabla f,\Delta f$ are uniformly bounded in $y$ (Assumption 5(i)), we conclude
@@
Proof of Corollary 2.
As in the proof of Theorem 3, we have
@@
Proof of Theorem 4.
The proof continues the footprint of Theorem 3. By definition, we have
@@
We deal with each term separately. First, by Lemma 1(ii), we have
@@
where $O(\eta^{3})$ is uniform in $y$ under Assumption 1 for up to $k=6$.
@@
Here the $O(\eta^{3})$ term is uniform in $y$ under Assumption 1 for up to $k=3$, and Assumption 5(i) & (ii).
For the determinant term, by Lemma 4, we have
@@
Again, the $O(\eta^{3})$ term is uniform in $y$ ensured by Assumption 5(i) & (ii).
@@
Recall that in the proof of Theorem 3, we have shown that $\mathcal{E}_{1}[q,f]=0$ uniformly in $y$ when $f=f_{*}$. Note also by Lemma 1(ii), we have
@@
We will first verify that $\mathcal{E}_{2}^{\mathrm{eff}}[q,f,g]=0$ when $f=f_{*}$ and $g=g_{*}$, and then use Lemma 1(ii) to conclude that $\mathcal{E}_{2}^{\mathrm{eff}}[p,f,g]-\mathcal{E}_{2}^{\mathrm{eff}}[q,f,g]=O(\eta)$ uniformly in $y$ to complete the proof
@@
where the last line uses Lemma 3 for the middle bracketed term. Recall that $f=f_{*}$ and $g=g_{*}$ satisfy
@@
using Lemma 2. Therefore, we conclude that $\mathcal{E}_{2}^{\mathrm{eff}}[q,f,g]=0$ when $f=f_{*}$ and $g=g_{*}$. ∎
@@ Appendix B Supporting Lemmas
Lemma 1.
Let $X\in\mathbb{R}^{d}$ be a random variable with density $p:\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$, and $Z\in\mathbb{R}^{d}$ be an random variable with density $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$. Consider the noisy measurement $Y=X+\sqrt{2\eta}Z$ with density $q:\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$.
(i) Let Assumption 2(i) holds, and Assumption 1 holds for up to $k=4$. Then the density $q$ admits the following expansion
@@
(ii) Let Assumptions 2(i) & (ii) hold, and let Assumption 1 holds for up to $k=6$, then
@@
Moreover, the derivatives of $q(y)$ admits the expansion
@@
[…]
where the last equality follows from Assumption 1 for $k=4$, with a constant depending only on $\|\nabla^{4}p\|_{\infty}$ and the fourth moment of $Z$. The $O(\eta^{2})$ is uniform over $y$ and $\eta\in[0,1)$. Recall $\langle\nabla^{2}p(y),\mathop{\mathbb{E}}[Z^{\otimes 2}]\rangle=\Delta p(y)$ by Assumption 2(i), the first part of the lemma is proved.
@@
[…]
where the last term is bounded by $C\cdot\eta^{3}$ for some constant $C$ depending only on $\|\nabla^{6}p\|_{\infty}$ and the sixth moment of $Z$. By Assumption 2(ii), we can verify that
@@
recall Lemma 2(ii), and (i), we have
@@
[…]
Proof of Lemma 4.
@@
[…]

### v4

@@ 5.1 First-Order Denoising Equation
Theorem 5.3 (Distributional Matching: Optimal First-Order Denoiser).
Under Assumptions 5.1(i) and 5.2(i), we consider the first-order denoiser $\mathbf{T}_{1}^{f}(y)$.
Let $p=P_{X}$ satisfy Assumption 3.1 for up to $k=4$, and the noise $Z$ has bounded 4-th moments as in Assumption 3.2(i). Then with some universal constant $C_{\Omega}>0$,
@@
An immediate corollary is a lower bound showing that the classical Bayes-optimal denoiser only achieves $\Theta(\sigma^{2})$ accuracy in terms of distributional matching, thus is suboptimal compared to the first-order denoiser $\mathbf{T}_{1}^{f_{*}}$, in the small constant noise regime.
Corollary 5.4 (Distributional Matching: Bayes-optimal Denoiser).
Consider the same setting and assumptions as in Theorem 5.3. The Bayes-optimal denoiser
@@ 5.2 Second-Order Denoising Equation
Theorem 5.5 (Distributional Matching: Optimal Second-Order Denoiser).
Under Assumptions 5.1(i) & (ii) and 5.2 (i) & (ii), we consider the second-order denoiser $\mathbf{T}_{2}^{f,g}$.
Let $p=P_{X}$ satisfy Assumption 3.1 for up to $k=6$, and the noise $Z$ has bounded 6-th moments as in Assumption 3.2(i) & (ii). Then with some universal constant $C_{\Omega}>0$,
@@
Again, this theorem requires stronger conditions than Theorem 5.3, in terms of smoothness of the signal distribution $P_{X}$ (up to 6-th order), and boundedness of 6-th moments and a Gaussian-like 4-th moment assumption, stated in Assumption 3.2(ii). If instead one wishes to consider general noise $Z$ with i.i.d. coordinates that have general 4-th moments $\mathop{\mathbb{E}}[Z_{i}^{4}]=\kappa\in[1,\infty)$, then the second-order denoising equation will be modified by adding a term of the form $(3-\kappa)\partial_{i}^{3}q$ for each coordinate $i$ in (5.2). We choose to state a less general version for simplicity of presentation.
@@ 6 Numerics
In both experiments, we observe the following phenomenon. First, the Bayes-optimal denoiser introduces an overly aggressive shrinkage, thus shrinking the distribution $\mathbf{T}^{\star}\sharp P_{Y}$ inside the mass of $P_{X}$ and producing poor distributional matching. In the mixture of Gaussian experiment, the $W(\mathbf{T}^{\star}\sharp P_{Y},P_{X})=0.214$, not a good improvement compared to no shrinkage $W(P_{Y},P_{X})=0.233$. In contrast, both the first-order and second-order denoisers produce significantly better distributional matching, with $W(\mathbf{T}_{1}^{f_{*}}\sharp P_{Y},P_{X})=0.050$ and $W(\mathbf{T}_{2}^{f_{*},g_{*}}\sharp P_{Y},P_{X})=0.038$, an order of magnitude improvement. Second, although visually the first-order and second-order denoisers look similar, the second-order denoiser achieves a better Wasserstein distance in both experiments by a margin, demonstrating the benefit of higher-order denoising. Both of our proposed denoisers significantly outperform the Bayes-optimal denoiser in terms of distributional matching, both visually and quantitatively. This is compatible with our theoretical findings in Theorems 4.3-5.5.
@@ A.1 Proofs for Section 4
$\mathop{\mathbb{E}}_{Y\sim P_{Y}}[m(\mathbf{T}_{1}^{f}(Y))]-\mathop{\mathbb{E}}_{X\sim P_{X}}[m(X)]$ $=\int_{\mathbb{R}^{d}}m(y+\eta\nabla f(y))q(y)\,\mathrm{d}y-\int_{\mathbb{R}^{d}}m(x)p(x)\,\mathrm{d}x\;.$ (A.1)
@@
$\qquad+\eta\int_{\mathbb{R}^{d}}\langle\nabla m(y),\nabla f(y)\rangle q(y)\,\mathrm{d}y$ (A.3)
$\qquad+\frac{\eta^{2}}{2}\int_{\mathbb{R}^{d}}\langle\nabla^{2}m(y+\delta\eta\nabla f(y)),\nabla f(y)\otimes\nabla f(y)\rangle q(y)\,\mathrm{d}y\;.$ (A.4)
@@
$\int_{\mathbb{R}^{d}}|\partial_{i}f(y)\partial_{j}f(y)|q(y)\,\mathrm{d}y\leq\left(\int_{\mathbb{R}^{d}}|\partial_{i}f(y)|^{2}q(y)\,\mathrm{d}y\right)^{1/2}\left(\int_{\mathbb{R}^{d}}|\partial_{j}f(y)|^{2}q(y)\,\mathrm{d}y\right)^{1/2}$
if (i) $\int_{\mathbb{R}^{d}}\|\nabla f(y)\|^{2}q(y)\,\mathrm{d}y<\infty$ and (ii) $\sup_{y\in\mathbb{R}^{d}}~\|\nabla^{2}m(y)\|_{\infty}<\infty$, we know the term in (A.4) is bounded by $O(\eta^{2})$. The constant here is independent of $y$, only depending on $m$.
Moving to the first term on the right-hand side of (A.2), by Lemma B.1(i), we have
@@
$\int_{\mathbb{R}^{d}}m(y)q(y)\,\mathrm{d}y-\int_{\mathbb{R}^{d}}m(y)p(y)\,\mathrm{d}y$ $=\eta\int_{\mathbb{R}^{d}}m(y)\Delta p(y)\,\mathrm{d}y+O(\eta^{2})\;,~\text{by (i)}$
$=-\eta\int_{\mathbb{R}^{d}}\langle\nabla m(y),\nabla p(y)\rangle\,\mathrm{d}y+O(\eta^{2})\;,~\text{by (ii)}$
$=-\eta\int_{\mathbb{R}^{d}}\langle\nabla m(y),\nabla q(y)\rangle\,\mathrm{d}y+O(\eta^{2})\;,~\text{by (i)}$
if (i) $\int_{\mathbb{R}^{d}}|m(y)|\,\mathrm{d}y<\infty$, $\int_{\mathbb{R}^{d}}|\partial_{i}m(y)|\,\mathrm{d}y<\infty$, and (ii) $m\nabla p$ vanishes as $y\rightarrow\infty$ because $m$ vanishes and $\nabla p$ is bounded due to Assumption 3.1. To be more specific, (ii) is a direct application of integration by parts
@@
and thus $\int_{\mathbb{R}^{d}}m\Delta p\,\mathrm{d}y=-\int_{\mathbb{R}^{d}}\langle\nabla m,\nabla p\rangle\,\mathrm{d}y$.
@@
$\mathop{\mathbb{E}}_{Y\sim P_{Y}}[m(\mathbf{T}_{1}^{f}(Y))]-\mathop{\mathbb{E}}_{X\sim P_{X}}[m(X)]$ $=\eta\int_{\mathbb{R}^{d}}\langle\nabla m(y),q(y)\nabla f(y)-\nabla q(y)\rangle\,\mathrm{d}y+O(\eta^{2})\;.$
@@
Proof of Corollary 4.4.
Recall the last equation in the proof of Theorem 4.3,
$\mathop{\mathbb{E}}_{Y\sim P_{Y}}[m(\mathbf{T}_{1}^{f}(Y))]-\mathop{\mathbb{E}}_{X\sim P_{X}}[m(X)]$ $=\eta\int_{\mathbb{R}^{d}}\langle\nabla m(y),q(y)\nabla f(y)-\nabla q(y)\rangle\,\mathrm{d}y+O(\eta^{2})\;.$
@@
Proof of Theorem 4.6.
Let $\eta:=\sigma^{2}/2$ to streamline the calculations. For simplicity, we use $f=f_{*}$ and $g=g_{*}$, defined in (3.2), in the following proof whenever there is no ambiguity.
@@
$\mathop{\mathbb{E}}_{Y\sim P_{Y}}[m(\mathbf{T}_{2}^{f,g}(Y))]-\mathop{\mathbb{E}}_{X\sim P_{X}}[m(X)]$ $=\int_{\mathbb{R}^{d}}m\left(y+\eta\nabla f(y)+\frac{\eta^{2}}{2}\nabla g(y)\right)q(y)\,\mathrm{d}y-\int_{\mathbb{R}^{d}}m(x)p(x)\,\mathrm{d}x\;.$
@@
The constant in the term $O(\eta^{3})$ depends on $\sup_{y\in\mathbb{R}^{d}}\|\nabla^{3}m(y)\|_{\infty}$ and $\|\nabla h(y)\|^{3}$, with the first bounded (Assumption 3.1) and second integrable (Assumption 4.1).
@@
$\int_{\mathbb{R}^{d}}m(y)q(y)\,\mathrm{d}y-\int_{\mathbb{R}^{d}}m(y)p(y)\,\mathrm{d}y$ $=\eta\int_{\mathbb{R}^{d}}m(y)\Delta p(y)\,\mathrm{d}y+\frac{\eta^{2}}{2}\int_{\mathbb{R}^{d}}m(y)\Delta^{2}p(y)\,\mathrm{d}y+O(\eta^{3})\;,$
$=-\eta\int_{\mathbb{R}^{d}}\langle\nabla m(y),\nabla p(y)\rangle\,\mathrm{d}y$
$\qquad-\frac{\eta^{2}}{2}\int_{\mathbb{R}^{d}}\langle\nabla m(y),\nabla\Delta p(y)\rangle\,\mathrm{d}y+O(\eta^{3})\;,$
@@
and the boundary condition as $m\nabla p,m\nabla\Delta p$ vanishes on the boundary as $y\rightarrow\infty$. Here we use the fact that $m$ vanishes and $\nabla p,\nabla\Delta p$ are bounded (Assumption 3.1). Now invoke Lemma B.1(ii) again, we have element-wise the following expansions
@@
$-\eta\int_{\mathbb{R}^{d}}\langle\nabla m(y),\nabla p(y)\rangle\,\mathrm{d}y-\frac{\eta^{2}}{2}\int_{\mathbb{R}^{d}}\langle\nabla m(y),\nabla\Delta p(y)\rangle\,\mathrm{d}y$
@@
where the two lines follows from $\int_{\mathbb{R}^{d}}|\partial_{i}m(y)|\,\mathrm{d}y<\infty$, for all $i$.
@@
Note that $q\langle\nabla m,\nabla f\rangle\nabla f$ vanishes on the boundary as $y\rightarrow\infty$ due to the fact that $\partial_{i}m$ vanishes and $q\partial_{i}f\partial_{j}f$ is bounded (Assumption 4.1). Therefore, applying integration by parts and recalling (A.7), we have
$\int_{\mathbb{R}^{d}}\langle\nabla^{2}m(y),\nabla f(y)\otimes\nabla f(y)\rangle q(y)\,\mathrm{d}y$
$=-\int_{\mathbb{R}^{d}}\Big[\langle\nabla m(y),\nabla^{2}f(y)\nabla f(y)\rangle q(y)+\langle\nabla m(y),\nabla f(y)\rangle\langle\nabla f(y),\nabla q(y)\rangle$
$\hskip 18.49988pt\hskip 18.49988pt\qquad+\langle\nabla m(y),\Delta f(y)\nabla f(y)\rangle q(y)\Big]\,\mathrm{d}y\;.$
@@
$=\frac{\eta^{2}}{2}\int_{\mathbb{R}^{d}}\Big\langle\nabla m,q\nabla g-(\nabla^{2}f\nabla f+\Delta f\nabla f)q-\langle\nabla f,\nabla q\rangle\nabla f+\nabla\Delta q\Big\rangle\,\mathrm{d}y+O(\eta^{3})\;.$
Using Lemma B.2, and recall that $f_{*}=\log q$, we have
@@ A.2 Proofs for Section 5
Proof of Theorem 5.3.
Recall the scaling $\eta:=\sigma^{2}/2$, and set $f=f^{\star}$ for simplicity of notation. By definition of $d_{\mathrm{MA}}$, and that $p=P_{X}$, $q=P_{Y}$, we have
@@
where $O(\eta^{2})$ is uniform in $y$ under Assumption 3.1 for up to $k=4$.
@@
Here the $O(\eta^{2})$ term is uniform in $y$ under Assumption 3.1 for up to $k=2$ and Assumption 5.2(i).
Moreover, by the expansion of the determinant, and Assumption 5.2(i), we have
@@
Again, the $O(\eta^{2})$ term is uniform in $y$ ensured by Assumption 5.2(i).
@@
Together with the assumption that $\nabla f,\Delta f$ are uniformly bounded in $y$ (Assumption 5.2(i)), we conclude
@@
Proof of Corollary 5.4.
As in the proof of Theorem 5.3, we have
@@
Proof of Theorem 5.5.
The proof continues the footprint of Theorem 5.3. By definition, we have
@@
We deal with each term separately. First, by Lemma B.1(ii), we have
@@
where $O(\eta^{3})$ is uniform in $y$ under Assumption 3.1 for up to $k=6$.
@@
Here the $O(\eta^{3})$ term is uniform in $y$ under Assumption 3.1 for up to $k=3$, and Assumption 5.2(i) & (ii).
For the determinant term, by Lemma B.4, we have
@@
Again, the $O(\eta^{3})$ term is uniform in $y$ ensured by Assumption 5.2(i) & (ii).
@@
Recall that in the proof of Theorem 5.3, we have shown that $\mathcal{E}_{1}[q,f]=0$ uniformly in $y$ when $f=f_{*}$. Note also by Lemma B.1(ii), we have
@@
We will first verify that $\mathcal{E}_{2}^{\mathrm{eff}}[q,f,g]=0$ when $f=f_{*}$ and $g=g_{*}$, and then use Lemma B.1(ii) to conclude that $\mathcal{E}_{2}^{\mathrm{eff}}[p,f,g]-\mathcal{E}_{2}^{\mathrm{eff}}[q,f,g]=O(\eta)$ uniformly in $y$ to complete the proof
@@
where the last line uses Lemma B.3 for the middle bracketed term. Recall that $f=f_{*}$ and $g=g_{*}$ satisfy
@@
using Lemma B.2. Therefore, we conclude that $\mathcal{E}_{2}^{\mathrm{eff}}[q,f,g]=0$ when $f=f_{*}$ and $g=g_{*}$. ∎
@@ Appendix B Supporting Lemmas
Lemma B.1.
Let $X\in\mathbb{R}^{d}$ be a random variable with density $p:\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$, and $Z\in\mathbb{R}^{d}$ be a random variable with density $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$. Consider the noisy measurement $Y=X+\sqrt{2\eta}Z$ with density $q:\mathbb{R}^{d}\rightarrow\mathbb{R}_{\geq 0}$.
(i) Let Assumption 3.2(i) hold, and Assumption 3.1 holds for up to $k=4$. Then the density $q$ admits the following expansion
@@
(ii) Let Assumptions 3.2(i) & (ii) hold, and let Assumption 3.1 hold for up to $k=6$, then
@@
Moreover, the derivatives of $q(y)$ admit the expansion
@@
$\left|q(y)-\left[p(y)+\frac{(\sqrt{2\eta})^{2}}{2!}\langle\nabla^{2}p(y),\mathop{\mathbb{E}}[Z^{\otimes 2}]\rangle\right]\right|$
$\qquad\leq\int_{\mathbb{R}^{d}}\frac{(\sqrt{2\eta})^{4}}{4!}\langle|\nabla^{4}p(y{+}\delta(-\sqrt{2\eta}z))|,|z^{\otimes 4}|\rangle\phi(z)\,\mathrm{d}z=C\cdot\eta^{2}\;,$
where the last equality follows from Assumption 3.1 for $k=4$, with a constant depending only on $\|\nabla^{4}p\|_{\infty}$ and the fourth moment of $Z$. The $O(\eta^{2})$ is uniform over $y$ and $\eta\in[0,1)$. Recall $\langle\nabla^{2}p(y),\mathop{\mathbb{E}}[Z^{\otimes 2}]\rangle=\Delta p(y)$ by Assumption 3.2(i), the first part of the lemma is proved.
@@
$\left|q(y)-\left[p(y)+\eta\Delta p(y)+\frac{(\sqrt{2\eta})^{4}}{4!}\langle\nabla^{4}p(y),\mathop{\mathbb{E}}[Z^{\otimes 4}]\rangle\right]\right|$
$\qquad\leq\int_{\mathbb{R}^{d}}\frac{(\sqrt{2\eta})^{6}}{6!}\langle|\nabla^{6}p(y{+}\delta(-\sqrt{2\eta}z))|,|z^{\otimes 6}|\rangle\phi(z)\,\mathrm{d}z\;,$
where the last term is bounded by $C\cdot\eta^{3}$ for some constant $C$ depending only on $\|\nabla^{6}p\|_{\infty}$ and the sixth moment of $Z$. By Assumption 3.2(ii), we can verify that
@@
recall Lemma B.2(ii), and (i), we have
@@
$\det(I_{d}+\eta A)=1+\eta\mathrm{tr}(A)+\frac{\eta^{2}}{2}\left(\mathrm{tr}(A)^{2}-\mathrm{tr}(A^{2})\right)+O(\eta^{3})\;.$
Proof of Lemma B.4.
@@
Noting that $\mathrm{tr}(A)=\sum_{i=1}^{d}\lambda_{i}$ and $\mathrm{tr}(A^{2})=\sum_{i=1}^{d}\lambda_{i}^{2}$, we complete the proof. ∎